A Complete Bibliography of Bell Labs Technical Journal, 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 October 2023 Version 1.01

Title word cross-reference

Augmented [KW20].

future [KW20].

Intelligence [KW20].

References

Kennedy:2020:FAI [KW20] S. Kennedy and C. White. The future

of augmented intelligence. Bell Labs Technical Journal, 25(??):1–18, 2020.

CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).